# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v4

@@ Abstract
The rapid advancements in artificial intelligence (AI), particularly in generative AI and large language models (LLMs), have profoundly impacted the creative industries, enabling more innovative content creation, enhancing workflows, and democratizing access to creative tools. This paper explores these technological shifts, with particular focus on how those that have emerged since our previous review in 2022 have expanded creative opportunities and improved efficiency. These technological advancements have enhanced the capabilities of text-to-image, text-to-video, and multimodal generation technologies. In particular, key breakthroughs in LLMs have established new benchmarks in conversational AI, while advancements in image generators have revolutionized content creation. We also discuss the integration of AI into post-production workflows, which has significantly accelerated and improved traditional processes. Once content has been created, it must be delivered to its audiences; the media industry is now facing the demands of increased communication traffic due to creative content. We therefore include a discussion of how AI is beginning to transform the way we represent and compress media content. We highlight the trend toward unified AI frameworks capable of addressing and integrating multiple creative tasks, and we underscore the importance of human insight to drive the creative process and oversight to mitigate AI-generated inaccuracies. Finally, we explore AI’s future potential in the creative sector, stressing the need to navigate emerging challenges and to maximize its benefits while addressing the associated risks.
Acknowledgements
@@ 1 Introduction
The influence of artificial intelligence (AI) has grown dramatically over the past few years, particularly due to the rise of generative AI and large language models (LLMs). These advancements are widely regarded as beneficial by many countries, creating significant opportunities for growth (e.g. as outlined in the UK, by the Authority of the House of Lords [1]). These advances have also had significant direct and indirect impacts on the creative industries, influencing the direction of their growth. Generative AI, for instance, primarily focuses on generating new data that is not identical to the training data yet shares similarities with it. However, the cardinality of the training data can be huge, larger than what any individual human has ever encountered. The resulting output may therefore act as a new source of inspiration.
@@
AI tools also provide opportunities for a wider range of users to work more efficiently and effectively, with even greater creativity. Moreover, these new technologies not only influence creators, but they also enable new ways for audiences to experience art and culture [2].
@@
A major breakthrough in generative AI has been led by OpenAI [^1] https://openai.com/, an AI research and deployment company, with their introduction of Generative Pre-trained Transformer (GPT) models for LLMs. LLMs are specifically designed to understand and generate human language. They are characterized by their vast size in terms of parameters and the amount of training data used to create them. This breakthrough was particularly impactful when the company released ChatGPT in 2022, which was fine-tuned from a model in the GPT-3.5 series. ChatGPT is a conversational model that includes advanced safety features that mitigate the generation of inappropriate content. Several other LLM platforms were also developed contemporaneously, such as LaMDA and PaLM by Google AI, Ernie Bot by Baidu, and BLOOM by BigScience. Additionally, Anthropic launched Claude, the LLM trained specifically to be harmless and honest, leveraging reinforcement learning from human feedback (RLHF) - a technique used to train AI systems to appear more human [3]. Nonetheless, ChatGPT stands out as the most renowned, thanks to its quick and efficient responses, and notably its public accessibility, being available for free.
@@
Another breakthrough in 2022 was in the area of text-to-image models. OpenAI achieved a significant milestone with DALL·E 2, producing impressive artworks and photorealistic images despite its limited language understanding. Midjourney by Midjourney, Inc., another well-known text-to-image generator, supports higher resolution images, up to 4096$\times$4096 pixels. Stable Diffusion by Stability AI, for which the code and model weights are publicly available [^2] https://github.com/Stability-AI/stablediffusion, allows developers and artists to further adapt AI to suit their own specific applications.
@@
The next breakthrough happened in 2023 when OpenAI unveiled GPT-4, a significantly larger model with estimated 1.8 trillion parameters and improved performance compared to its predecessors [4]. However, this still represents less than 1% of the human brain’s approximately 600 trillion synaptic connections [^3] https://www.rsb.org.uk/biologist-features/ai-versus-the-brain. GPT-4 is a multimodal large language model that can generate responses to both text and images. It incorporates DALL·E 3, enabling it to comprehend a much broader range of nuances and details than earlier versions. In March 2024, Claude 3 Opus by Anthropic was released, boasting multimodal capabilities in generating images, tables, graphs, and diagrams. Moreover, Anthropic claims that Claude 3 Opus outperforms GPT-4 in generating human-like dialog and contextually aware responses. These rapid advances have, in turn, led the creative industries to face significant challenges. For example, DMG Media, the Financial Times, and Guardian Media Group have highlighted concerns about the potential impact on print journalism, particularly if AI tools reduce the need for users to click through to news websites, affecting advertising and subscription revenues [1]. There is also concern about ‘AI-generated slop’—low-quality, mass-produced content created by AI that often lacks coherence or originality [^4] https://reutersinstitute.politics.ox.ac.uk/news/ai-generated-slop-quietly-conquering-internet-it-threat-journalism-or-problem-will-fix-itself. It is typically used for spam, search engines, or clickbait, and is criticized for cluttering the internet and undermining genuine human-created content.
@@
The generation of videos is significantly more challenging for AI than generating images. In February 2024, Google announced Gemini 1.5, which had the capability to process approximately 8 times more data than GPT-4, opening opportunities for video and audio processing [^5] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/. In the same month, OpenAI provided its first preview of Sora, a model capable of generating impressive realistic videos up to 1 minute long. Based on the videos released by OpenAI, Sora appears to outperform other text-to-video models. Sora is currently available to ChatGPT subscribers. A month later, Gemini 1.5 announced its support for native audio understanding in 180$+$ countries. With the emergence of these tools, together with the prospect of further advances, it is clear that video content creation will be a major beneficiary. This will further open up the media landscape for creativity and provide more opportunities for diverse storytellers, while also reducing production time. A recent example is the AI-generated Christmas commercial by Coca-Cola [^6] https://www.youtube.com/watch?v=4RSTupbfGog. Such advertisements overcome the limitations of current technologies by using very short videos with rapid scene transitions, ensuring that any artifacts, such as unnatural fingers, are less apparent.
@@
For the case of post-production workflows, generative AI may not have a direct impact, but the neural networks originally proposed for generative AI have been widely adapted to serve this purpose. This has led to significant improvements in both output quality and computational speed. Moreover, there is a noticeable trend towards adopting a unified framework rather than addressing individual tasks, as it better reflects real-world scenarios. For instance, natural history filmmaking involves challenging acquisition environments and high production standards. Filming often takes place in low light conditions, in the presence of heat haze, underwater or in adverse weather conditions. This often results in increased noise levels, focus issues, low contrast, color balance problems, and blurriness in the footage. In such cases, unified models can offer advantages in generalizing to diverse tasks and providing flexibility. Take Painter by BAAI Vision [5] as an example, which employs an image pair as a task prompt (similar to a text prompt in LLMs), their model transfers the input image to produce a similar output as the task prompt, enabling it to undertake various tasks such as segmentation, low-light enhancement or rain removal.
@@
One of the exciting aspects of using LLMs in the creative sector is that ‘The human in the loop’ [6] is simplified through text prompts, with sophisticated, multilingual language capabilities enabling artists to convey complex emotions and narratives. This is important because generative AI does produce mistakes, known as hallucinations. Human oversight is thus essential to correct this through reinforcement learning with feedback [7].
In this paper, the objective is to reveal to the reader the latest technology advancements that have emerged since our previous review paper on AI in the creative industries (published in 2022) [8]. Compared to this earlier paper, which was written when most AI technologies were used as support tools, this updated review describes the significant disruptive shifts that have emerged over the past 3-4 years driven by generative AI and other recent AI-based technologies. Similar to [8], we first provide a high-level overview of current advanced AI technologies (Section 2), followed by a selection of key creative domain applications (Section 3) where current AI technologies are changing creative practice. Finally, we discuss the challenges and the future potential of AI associated with the creative industries (Section 4).
@@ 2 Current Advanced AI Technologies
This paper provides a review of AI in the creative industries, building on our previous publication in 2022 [8]. The reader is referred to that work for an introduction to AI, basic neurons, convolutional neural networks (CNNs), generative adversarial networks (GANs), recurrent neural networks (RNNs) and deep reinforcement learning (DRL). In this paper, we emphasize four key technologies that have grown in importance since 2022 that have had a significant impact on the creative industries. These are Transformers, Large language models (LLMs), Diffusion Models (DMs), and Implicit Neural Representations (INRs). It is important to note that, while these newer technologies are gaining prominence, those from previous generations remain in widespread use, often in conjunction with the newer ones. For instance, CNNs complement transformers since CNNs effectively capture local features and semantic meaning, while the attention mechanism in transformers captures global dependencies.
One important class within AI that has become dominant since our previous review comprises Foundation models (FMs). These were described by The Stanford Institute for Human-Centered Artificial Intelligence in 2021 [9] as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. Foundation models have been enabled by rapid advances in AI-oriented computing power and have underpinned the emergence and success of Large Language Models, particularly following the launch of ChatGPT by OpenAI in 2022. ChatGPT has become the fastest-growing consumer software application in history [^7] https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/.
These technologies are expanded on below.
@@ 2.1 Transformers
In 2017, Google AI introduced the concept of ‘Transformer’ architectures in their publication ‘Attention Is All You Need’ [10]. This work has since been instrumental in the development and success of large language models alongside many other applications, including vision understanding [11], and multiple modality learning (e.g., Gato [12]).
@@
It should be noted that attention modules are not solely used in transformers, but have also been successfully integrated into other deep learning architectures such as CNNs, used for image classification [13], object detection [14], and other computer vision tasks [15].
In 2020, the first successful training of a transformer encoder for image recognition was published [11], referred to as a Vision Transformer (ViT). The ViT decomposes an input image into patches, similar to words in a sentence, and processes them through multi-head attention. Additionally, a Multilayer Perceptron (MLP) is employed as the feedforward network. In later work Microsoft introduced a hierarchical division of image inputs and a shifted window approach in their Swin Transformer [16]. This was reported to outperform ViT by 2.4% in ImageNet-22K classification (21,841 different categories). Its version 2 [17] applied a cosine function in the attention module. enabling the scaling of capacity and resolution. More detail on transformer-based object detection is discussed in Section 3.4.2. To date, Swin Transformers have been widely adopted in a range of applications including image restoration [18]. Comprehensive surveys on the use of transformers for image and video processing can be found in [19] and [20], respectively.
@@
In recent years, state space models [21, 22], commonly known as ‘Mamba’ have emerged. These are a linear variant of Transformers distinguished by their linear complexity in attention modeling. They are acknowledged to offer an equivalent or better performance than traditional Transformers, while demanding fewer computational resources and less memory.
Figure 1: Generative AI. (a) Transformer architecture [10]. (b) The top row represents the diffusion process and the bottom row represents the generation process of the new image [23]. (c) Latent Diffusion Models (LDM) [24].
@@ 2.2 Large language models
LLMs are based on transformer models using self-attention mechanisms as their core modules. Training LLMs comprises two steps: i) pre-training in an unsupervised learning manner, and ii) fine-tuning to a specific task or prompt-tuning for better user inputs. The models are first ‘pre-trained’ with a large amount of unlabelled text data to learn the meaning of words, and the relationships between those words, before using them to adapt to a downstream task. This is why OpenAI refers to their model as a Generative Pre-trained Transformer (GPT).
Fine-tuning involves training the model on new datasets. The drawback is, however, that these data need to be large enough to ensure generalization to new tasks. Prompt-tuning and prompt engineering are relatively new disciplines for developing and optimizing prompts to efficiently use language models. Prompts guide the way AI models interpret and respond to user queries. Prompt engineering is the process of structuring text or phrasing that guides the model towards generating the desired output. This relies heavily on trial and error, and an understanding of how the model responds. Prompt-tuning, on the other hand, involves training a small set of parameters before utilizing the LLM, thus requiring a relatively small amount of new data. This approach essentially converts text inputs into task-specific virtual inputs, referred to as tokens, while the pre-trained LLM remains unchanged [25]. The main drawback of prompt-tuning is the lack of interpretability. This paradigm has, however, extended to other domains, such as visual prompt tuning [26]. For a comprehensive survey of LLMs, please refer to [27].
To date, there are many LLM platforms as shown in Fig. 2. Fig. 2(a) shows their timeline. Many surveys and evaluations of LLMs are also available [27, 28, 29]. These include FLASK (Fine-grained Language Model Evaluation based on Alignment Skill Sets) [30] which evaluates LLMs based on 12 fine-grained skills for comprehensive language model evaluation: logical correctness, logical robustness, logical efficiency, factuality, commonsense understanding, comprehension, insightfulness, completeness, metacognition, conciseness, readability, and harmlessness. Evaluation results from FLASK are shown in Fig. 2 (b).
Figure 2: (a) Timeline of large language models. (b) Performance comparison evaluated by FLASK [30].

### v5

# Advances in Artificial Intelligence: A Review for the Creative Industries

@@ Abstract
Artificial intelligence (AI) has undergone transformative advances since 2022, particularly through generative AI, large language models (LLMs), and diffusion models, fundamentally reshaping the creative industries. However, existing reviews have not comprehensively addressed these recent breakthroughs and their integrated impact across the creative production pipeline. This paper addresses this gap by providing a systematic review of AI technologies that have emerged or matured since our 2022 review, examining their applications across content creation, information analysis, post-production enhancement, compression, and quality assessment. We document how transformers, LLMs, diffusion models, and implicit neural representations have established new capabilities in text-to-image/video generation, real-time 3D reconstruction, and unified multi-task frameworks—shifting AI from support tool to core creative technology. Beyond technological advances, we analyze the trend toward unified AI frameworks that integrate multiple creative tasks, replacing task-specific solutions. We critically examine the evolving role of human-AI collaboration, where human oversight remains essential for creative direction and mitigating AI hallucinations. Finally, we identify emerging challenges including copyright concerns, bias mitigation, computational demands, and the need for robust regulatory frameworks. This review provides researchers and practitioners with a comprehensive understanding of current AI capabilities, limitations, and future trajectories in creative applications.
keywords:
Creative industries, Machine learning, Deep learning, Generative AI, Large language models
@@ 1 Introduction
The influence of artificial intelligence (AI) has grown dramatically over the past few years, particularly due to the rise of generative AI and large language models (LLMs). These advancements are widely regarded as beneficial by many countries, creating significant opportunities for growth (e.g. as outlined in the UK, by the Authority of the House of Lords [UK:Large:2024]). These advances have also had significant direct and indirect impacts on the creative industries, influencing the direction of their growth. Generative AI, for instance, primarily focuses on generating new data that is not identical to the training data yet shares similarities with it. However, the cardinality of the training data can be huge, larger than what any individual human has ever encountered. The resulting output may therefore act as a new source of inspiration.
@@
AI tools also provide opportunities for a wider range of users to work more efficiently and effectively, with even greater creativity. Moreover, these new technologies not only influence creators, but they also enable new ways for audiences to experience art and culture [Jeary2024].
@@
A major breakthrough in generative AI has been led by OpenAI [^1] https://openai.com/, an AI research and deployment company, with their introduction of Generative Pre-trained Transformer (GPT) models for LLMs. LLMs are specifically designed to understand and generate human language. They are characterized by their vast size in terms of parameters and the amount of training data used to create them. This breakthrough was particularly impactful when the company released ChatGPT in 2022, which was fine-tuned from a model in the GPT-3.5 series. ChatGPT is a conversational model that includes advanced safety features that mitigate the generation of inappropriate content. Several other LLM platforms were also developed contemporaneously, such as LaMDA and PaLM by Google AI, Ernie Bot by Baidu, and BLOOM by BigScience. Additionally, Anthropic launched Claude, the LLM trained specifically to be harmless and honest, leveraging reinforcement learning from human feedback (RLHF) - a technique used to train AI systems to appear more human [Bai:Train:2021]. Nonetheless, ChatGPT stands out as the most renowned, thanks to its quick and efficient responses, and notably its public accessibility, being available for free.
@@
Another breakthrough in 2022 was in the area of text-to-image models. OpenAI achieved a significant milestone with DALL·E 2, producing impressive artworks and photorealistic images despite its limited language understanding. Midjourney by Midjourney, Inc., another well-known text-to-image generator, supports image resolutions of $\times1024$ pixels and also provides upscaling tools. Stable Diffusion by Stability AI, for which the code and model weights are publicly available [^2] https://github.com/Stability-AI/stablediffusion, allows developers and artists to further adapt AI to suit their own specific applications.
@@
The next breakthrough occurred in 2023 when OpenAI unveiled GPT-4, a significantly larger model with third-party estimates of 1.8 trillion parameters; OpenAI has not disclosed an official figure. It also demonstrated improved performance compared to its predecessors [openai:gpt4:2023]. However, this still many orders of magnitude smaller than the human brain’s estimated 100–1,000 trillion synaptic connections [^3] https://www.rsb.org.uk/biologist-features/ai-versus-the-brain. GPT-4 is a multimodal large language model that can generate responses to both text and images. In the ChatGPT interface, GPT-4 is integrated with DALL·E 3, enabling it to comprehend a much broader range of nuances and details than earlier versions. In March 2024, Claude 3 Opus by Anthropic was released, boasting multimodal capabilities in generating images, tables, graphs, and diagrams. Moreover, Anthropic claims that Claude 3 Opus outperforms GPT-4 in generating human-like dialog and contextually aware responses. These rapid advances have, in turn, led the creative industries to face significant challenges. For example, DMG Media, the Financial Times, and Guardian Media Group have highlighted concerns about the potential impact on print journalism, particularly if AI tools reduce the need for users to click through to news websites, affecting advertising and subscription revenues [UK:Large:2024]. There is also concern about ‘AI-generated slop’—low-quality, mass-produced content created by AI that often lacks coherence or originality [^4] https://reutersinstitute.politics.ox.ac.uk/news/ai-generated-slop-quietly-conquering-internet-it-threat-journalism-or-problem-will-fix-itself. It is typically used for spam, search engines, or clickbait, and is criticized for cluttering the internet and undermining genuine human-created content.
@@
The generation of videos is significantly more challenging for AI than generating images. In February 2024, Google announced Gemini 1.5, which had the capability to process approximately 8 times more data than GPT-4 ; this comparison refers to a 1M-token context window versus 128k tokens. This opening opportunities for video and audio processing [^5] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/. In the same month, OpenAI provided its first preview of Sora, a model capable of generating impressive realistic videos up to 1 minute long. Based on the videos released by OpenAI, Sora appears to outperform other text-to-video models. Sora is currently available to ChatGPT subscribers , but new Plus accounts may face a wait-list or monthly credit limits. A month later, Gemini 1.5 announced its support for native audio understanding in 180++ countries. With the emergence of these tools, together with the prospect of further advances, it is clear that video content creation will be a major beneficiary. This will further open up the media landscape for creativity and provide more opportunities for diverse storytellers, while also reducing production time. A recent example is the AI-generated Christmas commercial by Coca-Cola [^6] https://www.youtube.com/watch?v=4RSTupbfGog , which elicited mixed responses in the press [Yang2025], thereby pre-empting questions regarding potential hype. Such advertisements overcome the limitations of current technologies by using very short videos with rapid scene transitions, ensuring that any artifacts, such as unnatural fingers, are less apparent.
@@
For the case of post-production workflows, generative AI may not have a direct impact, but the neural networks originally proposed for generative AI have been widely adapted to serve this purpose. This has led to significant improvements in both output quality and computational speed. Moreover, there is a noticeable trend towards adopting a unified framework rather than addressing individual tasks, as it better reflects real-world scenarios. For instance, natural history filmmaking involves challenging acquisition environments and high production standards. Filming often takes place in low light conditions, in the presence of heat haze, underwater or in adverse weather conditions. This often results in increased noise levels, focus issues, low contrast, color balance problems, and blurriness in the footage. In such cases, unified models can offer advantages in generalizing to diverse tasks and providing flexibility. Take Painter by BAAI Vision [Wang:Painter:2023] as an example, which employs an image pair as a task prompt (similar to a text prompt in LLMs), their model transfers the input image to produce a similar output as the task prompt, enabling it to undertake various tasks such as segmentation, low-light enhancement or rain removal.
@@
One of the exciting aspects of using LLMs in the creative sector is that ‘The human in the loop’ [chung:human-loop:2021] is simplified through text prompts, with sophisticated, multilingual language capabilities enabling artists to convey complex emotions and narratives. This is important because generative AI does produce mistakes, known as hallucinations. Human oversight is thus essential to correct this through reinforcement learning with feedback [Wu:brief:2023].
In this paper, the objective is to present the major technological advancements that have emerged since our previous review on AI in the creative industries (2022) [Anantrasirichai:AI:2022]. Whereas the earlier paper was written at a time when AI tools primarily served supportive roles, this updated review captures the disruptive shifts driven by generative AI and related technologies over the past few years. Adopting an integrative narrative rather than a systematic approach, we aim to synthesize major developments and identify cross-domain themes, rather than conduct an exhaustive quantitative survey. Sources published between 2022 and mid-2025 were selected from leading AI conferences (CVPR, ICCV, ECCV, NeurIPS, ICLR, IEEE TPAMI, IEEE TIP), verified arXiv preprints highlighting emerging developments, and industry releases, based on their technical significance and impact (e.g., OpenAI, Google DeepMind, Meta, Adobe). Selection was guided by recency, originality, and relevance to creative-industry applications, ensuring representative coverage of key models and trends. In summary, we initially screened more than 1,000 documents published between 2022 and mid-2025, including major AI conference papers ($\approx$60%), journal articles ($\approx$30%), and industry or policy reports ($\approx$10%). From these, around 450 sources were retained for detailed analysis across creation, enhancement, immersive media, and AI-driven production workflows. The review first outlines recent advances in AI technologies (Section 2), then highlights their applications in creative domains (Section 3), and finally discusses emerging challenges and future directions (Section 4).
@@ 2 Current Advanced AI Technologies
This paper provides a review of AI in the creative industries, building on our previous publication in 2022 [Anantrasirichai:AI:2022]. The reader is referred to that work for an introduction to AI, basic neurons, convolutional neural networks (CNNs), generative adversarial networks (GANs), recurrent neural networks (RNNs) and deep reinforcement learning (DRL). In this paper, we emphasize four key technologies that have grown in importance since 2022 that have had a significant impact on the creative industries. These are Transformers, Large language models (LLMs), Diffusion Models (DMs), and Implicit Neural Representations (INRs). It is important to note that, while these newer technologies are gaining prominence, those from previous generations remain in widespread use, often in conjunction with the newer ones. For instance, CNNs complement transformers since CNNs effectively capture local features and semantic meaning, while the attention mechanism in transformers captures global dependencies.
One important class within AI that has become dominant since our previous review comprises Foundation models (FMs). These were described by The Stanford Institute for Human-Centered Artificial Intelligence in 2021 [Bommasani2021FoundationModels] as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. Foundation models have been enabled by rapid advances in AI-oriented computing power and have underpinned the emergence and success of Large Language Models, particularly following the launch of ChatGPT by OpenAI in 2022. ChatGPT has become the fastest-growing consumer software application in history [^7] https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/. These technologies are expanded on below.
@@ 2.1 Transformers
In 2017, Google AI introduced the concept of ‘Transformer’ architectures in their publication ‘Attention Is All You Need’ [Vaswani:attention:2017]. This work has since been instrumental in the development and success of large language models alongside many other applications, including vision understanding [Dosovitskiy:image:2021], and multiple modality learning (e.g., Gato [Reed:Generalist:2022]).
@@
It should be noted that attention modules are not solely used in transformers, but have also been successfully integrated into other deep learning architectures such as CNNs, used for image classification [Li:HAM:2022], object detection [Woo:CBAM:2018], and other computer vision tasks [Guo:Attention:2022].
In 2020, the first successful training of a transformer encoder for image recognition was published [Dosovitskiy:image:2021], referred to as a Vision Transformer (ViT). The ViT decomposes an input image into patches, similar to words in a sentence, and processes them through multi-head attention. Additionally, a Multilayer Perceptron (MLP) is employed as the feedforward network. In later work Microsoft introduced a hierarchical division of image inputs and a shifted window approach in their Swin Transformer [Liu:Swin:2021]. This was reported to outperform ViT by 2.4% in ImageNet-22K classification (21,841 different categories). Its version 2 [Liu:Swinv2:2022] applied a cosine function in the attention module. enabling the scaling of capacity and resolution. More detail on transformer-based object detection is discussed in Section 3.4.2. To date, Swin Transformers have been widely adopted in a range of applications including image restoration [Fan:SUNet:2022]. Comprehensive surveys on the use of transformers for image and video processing can be found in [Khan:Transformers:2022] and [Selva:video:2023], respectively.
@@
In recent years, state space models [gu2023mamba, zhu2024vision], commonly known as ‘Mamba’ have emerged. These are a linear variant of Transformers distinguished by their linear complexity in attention modeling. They are acknowledged to offer an equivalent or better performance than traditional Transformers, while demanding fewer computational resources and less memory.
Figure 1: Generative AI. (a) Transformer architecture [Vaswani:attention:2017]. (b) The top row represents the diffusion process and the bottom row represents the generation process of the new image [Yang:diffusion:2023]. (c) Latent Diffusion Models (LDM) [Rombach:LDM:2022].
@@ 2.2 Large language models
LLMs are based on transformer models using self-attention mechanisms as their core modules. Training comprises two steps: i) pre-training with large amounts of unlabelled text data in an unsupervised manner to learn word meanings and relationships, and ii) task adaptation through either fine-tuning or prompt-tuning. This pre-training approach is why OpenAI refers to their model as a Generative Pre-trained Transformer (GPT).
Fine-tuning involves training the model on new datasets, which must be large enough to ensure generalization to new tasks. Prompt-tuning and prompt engineering are emerging disciplines focused on developing and optimizing prompts for efficient model use. Prompts guide how AI models interpret and respond to user queries. Prompt engineering structures text or phrasing to steer the model toward desired outputs, relying heavily on experimentation and understanding model behavior. Prompt-tuning instead trains a small set of parameters before using the LLM, requiring relatively little new data. This approach converts text inputs into task-specific virtual tokens while keeping the pre-trained model unchanged [Lester:power:2021]. Its main drawback is reduced interpretability, though the paradigm has extended to other domains such as visual prompt tuning [Jia:VPT:2022]. For a comprehensive survey of LLMs, see [zhao:survey:2023].
To date, there are many LLM platforms as shown in Fig. 2. Fig. 2(a) shows their timeline. Many surveys and evaluations of LLMs are also available [zhao:survey:2023, Chang:Survey:2024, YAO:Survey:2024]. These include FLASK (Fine-grained Language Model Evaluation based on Alignment Skill Sets) [Ye:FLASK:2024] which evaluates LLMs based on 12 fine-grained skills for comprehensive language model evaluation: logical correctness, logical robustness, logical efficiency, factuality, commonsense understanding, comprehension, insightfulness, completeness, metacognition, conciseness, readability, and harmlessness. Evaluation results from FLASK are shown in Fig. 2 (b).
Figure 2: (a) Timeline of large language models. (b) Performance comparison evaluated by FLASK [Ye:FLASK:2024].
A clear divide exists between open-source and proprietary LLMs. Open-source models such as LLaMA 3, Mistral, and Falcon promote transparency, reproducibility, and creative customization, allowing researchers to fine-tune or deploy them locally with improved data control. Proprietary systems like GPT-4, Gemini 1.5, and Claude 3, meanwhile, deliver stronger multimodal integration, safety alignment, and reliability—benefiting from access to large, private datasets and compute resources. Yet their closed architectures limit interpretability and independent evaluation. Together, these ecosystems form a complementary landscape balancing openness and innovation against performance and safety, shaping accessibility and creative autonomy in AI-driven practice.